# ATRNet-STAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

### v3

@@ Abstract
Remote sensing, transformer, datasets and benchmarks, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model1 Introduction
@@
Synthetic Aperture Radar (SAR) imaging is capable of generating high resolution imagery irrespective of lighting conditions and weather [1, 2, 3, 4], and has become an indispensable tool for Earth Observation [5, 2, 6]. SAR Remote Sensing (RS) data enables analysis and recognition of objects and scenes [2, 3, 7], which has become a valuable complement to other RS imagery [8, 9, 10]. Consequently, as a fundamental and challenging field in RS image analysis, SAR automatic target recognition (ATR) [11, 12], which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings), has become an active research area for several decades (Fig. 2 (a)). SAR ATR has a wide range of civilian and military applications, including military reconnaissance, urban management, disaster assessment, emergency rescue, and global surveillance [13, 14, 15, 16]. Despite the remarkable achievement achieved over the past several decades in the field of SAR ATR, the accurate, robust, and efficient recognition of any target in an open world remains unresolved [11, 7].
@@
Need for ATRNet The advent of the big data era has propelled the evolution of RS pretraining foundation models [17, 18, 19, 20] where large-scale pre-training enables efficient cross-task adaptation with minimal finetuning. However, the scarcity of large-scale standardized datasets restricts the development of generalizable data-driven SAR ATR methods, compared to the success of foundation models in other RS sensors [21, 22, 23]. The data sensitivity, acquisition costs, annotation difficulty, and complexity of SAR imaging hinder the establishment of open large-scale data ecosystems: More than 50% of SAR ATR studies still rely on the 1990s-era Moving and Stationary Target Acquisition and Recognition (MSTAR) [24] dataset (Fig. 2 (b)). In addition, non-standardized evaluation protocols across existing benchmarks (e.g., MSTAR, OpenSARShip [25], and FUSAR-Ship [26]) impede objective algorithm comparison. Therefore, building a large SAR ATR dataset and benchmark is necessary to unlock new model capabilities in this field. We aim for ATRNet—a massive, diversified SAR target dataset benchmark for modeling and recognizing target characteristics.
@@
Need for ATRNet-STAR. Researchers have struggled to construct many SAR target datasets to study SAR ATR algorithms. Many SAR classification datasets in Table II have significantly improved data diversity. In particular, new SAR detection datasets [27, 28] have emerged with 100,000 images. However, our previous research [29, 30] on SAR foundation models revealed that collecting publicly available datasets yields fewer than 300,000 target samples, with severe sample imbalance: mainly ship detection datasets. As the inaugural phase of ATRNet, we focus on vehicle targets due to: (1) Recent datasets predominantly are targets with spaceborne SAR and constrained imaging geometries, whereas airborne platforms offer higher resolution (0.1-0.3m) and richer imaging conditions. (2) Vehicle recognition has driven SAR ATR research for three decades since the building of MSTAR, establishing mature research systems. (3) The 1990s-era MSTAR dataset, despite its seminal role, suffers from idealized acquisition conditions and saturated performance (near 99% accuracy), failing to reflect real-world complexities and support innovation in the 2020s.
@@
| Location | ideally centered | random, noncentered |
@@
Our Objective and Solution. As a first step towards building a large-scale SAR target dataset, we introduce ATRNet-STAR111STAR: Stationary Target Acquisition and Recognition. This dataset focuses on the fine-grained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task in different condictions., a fine-grained dataset for vehicle target recognition as a new breakthrough to the previous benchmark dataset MSTAR in Table. I. We completed program designation, data acquisition and processing, and benchmark construction with nearly two years of effort and considerable resources. In particular, ATRNet-STAR collects 194,324 object images from 40 target types, 5 scenes, and various imaging conditions with detailed annotation. It is the largest public SAR vehicle recognition dataset, even 10 times the size of any similar vehicle dataset previously available, as shown in Table. II. Its sufficient target samples can support multiple aspects of generation, detection, and classification research. To facilitate innovation and comparisons, we built a well-designed classification and detection benchmark, including 15 representative methods based on 7 experimental settings for robust ATR, few-shot, and transfer learning. Our results show that SAR ATR under complex conditions is still very difficult, and representative methods that performed well in previous datasets can not solve some of our difficult experimental conditions, such as image angle and scene variations. We also find the effectiveness of the pre-trained foundation model compared to other methods, which illustrates the importance of a large-scale dataset. Besides, the transfer learning results of various models pre-trained on this dataset to other datasets show that our dataset benefits the feature learning of different ground targets. Finally, We provide detailed analyses and discussions to explore important issues for future research. We believe that our new dataset, with comprehensive imaging conditions and experiment benchmark, will further promote the promise and reproducible research in SAR ATR.
Our Contributions. The main contributions of this paper are listed below.
@@
Dataset and Benchmarks. To our best knowledge, the ATRNet-STAR is the largest dataset towards SAR vehicle recognition in the wild. Building ATRNet-STAR is a challenging task, and has taken nearly two years. We derive benchmark datasets from ATRNet-STAR to facilitate the performance evaluation of target classification and detection methods.
@@
Code library. To the best of our knowledge, we are the first to integrate publicly available and representative methods into a single code library to facilitate reproducible and large-scale performance evaluations in this field.
@@
Performance Evaluation. We review most major approaches and provide a comparative study of 15 representative methods with 7 experimental settings on classification and detection benchmarks derived from ATRNet-STAR, offering official baseline references when developing new approaches.
@@
Insights and Future Directions. By analyzing the quantitative results, we provide insightful discussion that can guide the development of SAR ATR algorithms and provide potential future research directions in this field.
@@
This work mainly focuses on the introduction, performance evaluation, and potential value of ATRNet-STAR. The full annotated dataset, code library, and reported results are available at https://github.com/waterdisappear/ATRNet-STAR after acceptance. The remainder of this paper is organized as follows. Sec. 2 discusses related work in SAR target classification datasets and algorithms. Sec. 3 introduces our dataset ATRNet-STAR. Secs. 4 conduct extensive experiments to build the benchmark. Sec. 5 concludes the paper and discusses future work.
@@ 2 Related Work
Here, we review the datasets and algorithms for SAR target classification and finally discuss the significance of the datasets for algorithmic studies.
@@ 2.1 Datasets for SAR target classification
Datasets are the cornerstone of the development of new ATR techniques, especially in an era based on deep learning. In recent years, there has been notable progress in RS datasets across various modalities [31], including visible light [32, 33, 34, 35, 36], SAR [27, 28], infrared [37], and multimodal [38, 39]. However, not all sensors possess sufficient target samples to fully leverage the potential of big data and foundation models. In the past decades, the MSTAR [24] dataset has played a pivotal role in SAR target classification tasks. However, current deep learning-based methods are saturated on this dataset benchmark setting. MSTAR’s insufficient samples also suffers from data biases such as background correlation [11, 40, 41]. Therefore, many datasets been developed to promote the advancement of this field, especially in recent years. We analyze SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig. 3.
@@
| ATRNet-STAR | 2024 | airborne | X, Ku | quad | 21 | 40 | 0.12 0.15 | 128 | 194,324 | A large-scale fine-grained dataset |
@@
Vehicle datasets. Vehicle datasets typically include a wide variety of fine-grained vehicle types and diverse imaging conditions, designed to evaluate robust classification performance under varying training and test set. These SAR vehicle datasets are summarized in Fig. 5, which provides an overview of the target categories. MSTAR was released by the Defense Advanced Research Projects Agency (DARPA) as the first public dataset for SAR target classification. It contains SAR image slices of 10 military vehicles and a reference target under different imaging angles with spotlight SAR. However, MSTAR has a background correlation problem and ideal collection conditions. Compared to MSTAR, the QinetiQ [42] dataset provided a non-idealized acquisition condition with a non-centered target location and stronger background clutter. CV Domes [43] and Gotcha [44] constructed simulation and measurement samples of civilian vehicles. These two datasets are in the raw echo data format and can be used for SAR imaging research, but they are not as convenient for researchers to use as MSTAR. Besides, these two civilian vehicle datasets are not rich in target classes and acquisition conditions. SARSIM [45] and SAMPLE [48] are two simulation datasets designed to study transfer learning. In particular, SARSIM uses 14 CAD models, and SAMPLE has simulated and measured sample pairs based on partially MSTAR’s public vehicle. The simulation data have discrepancies with the measured data in background clutter, target scattering characteristics, and speckle noise. Although MSTAR includes 3 grass locations, most public images are from the same location. MGTD [47] collects the training and test sets in two different laboratory backgrounds. Another issue is that the depression angles of MSTAR public images are mostly in 15∘ and 17∘, with some classes missing at other depression angles. Hence, NUAAminiSAR [50] collects images of military targets at different depression angles by circular SAR in a national defense park, including 8 military vehicles and an airplane. As shown above, existing vehicle datasets lack sufficient diversity in target types, scene settings, and sensor conditions.
@@
Ship datasets. SAR ship datasets have emerged with the development of satellite SAR technology. The difference with the vehicle dataset is that the ship classification dataset usually uses a satellite platform, and the imaging angles are limited. Ship datasets can be constructed and annotated using global ship tracking systems, but label noise remains a significant challenge. OpenSARShip [25], HR4S [46], and FUSAR-Ship [26] advance the development of ship classification methods. These datasets are derived from different satellites and face difficult harbor backgrounds and class imbalance challenges. OpenSARShip collects medium-resolution images of 16 classes in harbors with Sentinel-1. HR4S constructs high-resolution ship images in ports and coastal areas with GF-3 and RADARSAT-2. FUSAR-Ship contains high-resolution images under various sea backgrounds with 15 ship main classes and 98 subclasses based on GF-3. However, OpenSARShip and FUSAR-Ship have a serious class imbalance and long-tailed distribution challenges. Researchers construct various experimental settings and use some target categories for the classification based on these two datasets.
@@
Aircraft datasets. Aircraft datasets have emerged in recent years, presenting unique challenges in RS. Due to their streamlined surfaces and less obvious scattering characteristics, aircraft are more difficult to classify than vehicle and ship targets. Multiangle Aircraft Target Dataset (MATD) [51] collects SAR images of two aircraft at different imaging angles with a drone strimap SAR equipment. SAR-ACD [52] contains 2 classes and 6 types of civil aircraft in 3 airports.
@@
Others. IRIS-SAR [49] is rich in aircraft, ships, and vehicle simulation types. It considers noise, clutter, and shadows in SAR image simulation. However, it does not take into account occlusion in complicated backgrounds.
@@
Object detection dataset. Beyond the classification datasets focused on individual targets, numerous SAR object detection datasets are designed to detect multiple targets. These datasets typically feature large image sizes, significant background interference, and complex targeting challenges. Examples include SARDet-100K [27] and FAIR-CSAR [28], mainly satellite-based datasets containing approximately 100,000 images with resolutions exceeding 1 meter. This resolution makes it hard to detect and classify vehicle targets. Their non-cooperation objectives make it difficult to label the target types and accurately control the characteristics of the target under different angles and scenes. Therefore, our dataset provides a more fine-grained SAR vehicle characterization dataset under target class and imaging conditions.
@@
Collecting target samples under different imaging conditions remains a challenging and resource-intensive endeavor. After thirty years, MSTAR remains the most used classification dataset. The evolution of SAR ATR necessitates a new large-scale dataset and benchmark in the new era. After the above investigation, we suggest that a good SAR ATR dataset should pay attention to the following key points.
@@
Richness. A comprehensive dataset should include diverse target categories, scenes, and imaging conditions to reflect real-world environments and reduce biases. It also needs to contain enough formats and metadata, such as complex data and imaging parameters, to help researchers in algorithm innovation.
@@
Standardization. An official dataset and benchmark are beneficial to fair research and comparisons. In addition, the dataset format and target system should be standardized to be compatible with mainstream datasets and for further expansion and scaling.
@@
Usability. The dataset should be readily accessible and user-friendly for researchers to download and utilize. Some datasets may be difficult to access or require additional SAR imaging and pre-processing steps.
@@
Therefore, we build and open-source a new large-scale dataset, ATRNet-STAR, with various categories and imaging conditions. This dataset is carefully labeled with extensive metadata and has an official benchmark based on 7 experimental settings and 15 comparison methods with 2 data formats. This dataset is our first step in building a new SAR target dataset system. In the future, we will expand the objects from vehicles to a wider category list to further promote SAR ATR research.

### v4

@@ Abstract
Remote sensing, datasets and benchmarks, performance evaluation, synthetic aperture radar, automatic target recognition, image classification, object detection, foundation model, transformer1 Introduction
@@
Synthetic Aperture Radar (SAR) imaging is capable of generating high-resolution imagery irrespective of lighting conditions and weather [1, 2, 3, 4], and has become an indispensable tool for Earth observation [5, 2, 6]. SAR Remote Sensing (RS) data enables analysis and recognition of objects and scenes [2, 3, 7], which has become a valuable complement to other RS imagery [8, 9, 10]. Consequently, as a fundamental and challenging field in RS image analysis, SAR automatic target recognition (ATR) [11, 12], which autonomously detects and classifies objects of interest (e.g., , vehicles, ships, aircraft, and buildings), has become an active research area for several decades (Fig. 2 (a)). SAR ATR has a wide range of civilian and military applications, including global surveillance, military reconnaissance, urban management, disaster assessment, and emergency rescue [13, 14, 15, 16, 17]. Despite the remarkable achievements over the past several decades in the field of SAR ATR, the accurate, robust, and efficient recognition of any target in an open world remains unresolved [11, 7].
@@
Need for ATRNet - The advent of big data has propelled the evolution of RS pre-training foundation models [18, 19, 20, 21] where large-scale pre-training enables efficient cross-task adaptation with minimal finetuning. However, the scarcity of large-scale standardized datasets restricts the development of generalizable data-driven SAR ATR methods, compared to the success of foundation models in other RS sensors [22, 23, 24]. The data sensitivity, acquisition costs, annotation difficulty, and complexity of SAR imaging hinder the establishment of open large-scale data ecosystems: More than 50% of SAR ATR studies still rely on the 1990s-era Moving and Stationary Target Acquisition and Recognition (MSTAR) [25] dataset (Fig. 2 (b)). In addition, non-standardized evaluation protocols across existing benchmarks (e.g., MSTAR, OpenSARShip [26], and FUSAR-Ship [27]) impede objective algorithm comparison. Therefore, building a large SAR ATR dataset and benchmark is necessary to unlock new model capabilities in this field. We aim for ATRNet—a massive, diverse, and standard SAR target dataset benchmark for model and recognize target characteristics.
@@
Need for ATRNet-STAR111STAR: Stationary Target Acquisition and Recognition. Our dataset focuses on the fine-grained single target recognition problem in SAR ATR, i.e., the image classification task, but we provide a horizontal bounding box label for the object detection task in different conditions. - Researchers have struggled to construct many SAR target datasets for ATR tasks. Many SAR classification datasets in Table II have significantly improved data diversity. In particular, new SAR detection datasets [28, 29] have emerged with 100,000 images. However, our previous research [30, 31] on SAR foundation models revealed that collecting public datasets yields fewer than 200,000 available target samples due to severe sample imbalance—mainly ship detection datasets. As the inaugural phase of ATRNet, we focus on vehicle targets due to: (1) Recent datasets predominantly are targets based on spaceborne SAR with constrained imaging geometries, whereas airborne platforms offer higher resolution (0.1-0.3m) and flexible imaging conditions in complex scenes. (2) Vehicle recognition has driven SAR ATR research for three decades since MSTAR in the 1990s, establishing mature research systems. (3) The 1990s-era MSTAR dataset, despite its seminal role, suffers from idealized acquisition conditions and saturated performance (near 99% accuracy), failing to reflect real-world complexities and support innovation in the 2020s.
@@
| Location | ideally centered | random, non-centered |
@@
Our Objective and Solution - As a first step towards building a large-scale SAR target dataset, we introduce ATRNet-STAR, a fine-grained dataset for vehicle target recognition as a new breakthrough over the previous benchmark dataset MSTAR in Table I. We completed program designation, data acquisition and processing, and benchmark construction with nearly two years of effort and considerable resources. In particular, ATRNet-STAR collects 194,324 object images from 40 target types, 5 scenes, and various imaging conditions with detailed annotation. It is the largest public SAR vehicle recognition dataset, even 10 times the size of any previous vehicle dataset, as shown in Table II. Its sufficient target samples can support various aspects of generation, detection, and classification research. To facilitate innovation and comparisons, we built a well-designed classification and detection benchmark, including 15 representative methods based on 7 experimental settings for robust ATR, few-shot, and transfer learning. Our results show that SAR ATR under complex conditions is still difficult, and representative methods that performed well in previous datasets cannot solve our difficult conditions, such as image angle and scene variations. We also find the effectiveness of the foundation model compared to others, which illustrates the importance of a large-scale dataset. Besides, the transfer learning results of various models pre-trained on this dataset to other datasets show that our dataset contributes to the feature representation of different ground targets. Finally, we provide detailed analyses and discussions to explore important issues for future research. We believe that our new dataset, with comprehensive imaging conditions and experiment benchmarks, will further promote the promise and reproducible research in SAR ATR.
Our Contributions - The main contributions of this paper are listed below.
@@
Dataset and Benchmarks - To the best of our knowledge, the ATRNet-STAR is the largest dataset for SAR vehicle recognition. Building ATRNet-STAR is a challenging task and has taken nearly two years. We derive benchmark datasets from ATRNet-STAR to facilitate the research innovation and performance evaluation of target classification and detection methods.
@@
Code library - It is worth pointing out that we are the first to integrate publicly available and representative methods into a single code library to facilitate reproducible and large-scale performance evaluations in this field.
@@
Performance Evaluation - We review most major approaches and provide a comparative study of 15 representative methods with 7 experimental settings on classification and detection benchmarks, named ATRBench, offering official baseline references when developing new approaches.
@@
Insights and Future Directions - By analyzing the quantitative results, we provide insightful discussion that can guide the development of SAR ATR algorithms and provide potential future research directions in this field.
@@
This work mainly focuses on the introduction, performance evaluation, and potential value of ATRNet-STAR. The annotated dataset, code library, and reported results are available at https://github.com/waterdisappear/ATRNet-STAR after acceptance.
@@ 2 Related Work
Here, we review the datasets and algorithms for SAR target classification and finally discuss the significance of the datasets.
@@ 2.1 Datasets for SAR target classification
Datasets are the cornerstone of the development of new ATR techniques, especially in the era of deep learning. In recent years, there has been notable progress in RS datasets in various modalities [32], including visible light [33, 34, 35, 36, 37], SAR [28, 29], infrared [38], and multimodal [39, 40]. However, not all modalities possess sufficient target samples to fully leverage the potential of big data and foundation models. In the past decades, the MSTAR [25] dataset has played a key role in SAR target classification tasks. However, current deep learning-based methods are saturated in this benchmark dataset setting. Insufficient samples in MSTAR also have data biases such as background correlation [11, 41, 42]. Therefore, many datasets have been developed to promote the advancement of this field, especially in recent years. We analyze SAR target classification datasets that focus on individual target characteristics from the 1990s to the 2020s, which are detailed listed in Table II and Fig. 3.
@@
| ATRNet-STAR | 2025 | airborne | X, Ku | quad | 21 | 40 | 0.12 0.15 | 128 | 194,324 | A large-scale fine-grained dataset |
@@
Vehicle datasets - Vehicle datasets usually have various fine-grained vehicle types and diverse imaging conditions designed to evaluate robust classification performance under varying training and test sets. These SAR vehicle datasets are summarized in Fig. 5, which provides an overview of the target classes. MSTAR was released by the Defense Advanced Research Projects Agency (DARPA) as the first public dataset for SAR target classification. It contains SAR image slices of 10 military vehicles and a reference target under different imaging angles with spotlight SAR. However, MSTAR has a background correlation problem and ideal collection conditions. Compared to MSTAR, the QinetiQ [43] dataset provided a non-idealized acquisition condition with a non-centered target location and stronger background clutter. CV Domes [44] and Gotcha [45] constructed simulation and measurement samples of civilian vehicles. These two datasets are in the raw echo data format and can be used for SAR imaging research, but they are not as convenient for researchers to use as MSTAR. Besides, these two civilian vehicle datasets are not rich in target classes and acquisition conditions. SARSIM [46] and SAMPLE [49] are two simulation datasets designed to study transfer learning. In particular, SARSIM uses 14 CAD models, and SAMPLE has simulated and measured sample pairs based on partially MSTAR’s vehicle. The simulation data have discrepancies with the measured data in background clutter, target scattering characteristics, and speckle noise. Although MSTAR includes 3 grasslands, most public images are from the same location. MGTD [48] collects training and test sets in two different laboratory backgrounds. Another issue is that the depression angles of the MSTAR public images are mostly in 15∘ and 17∘, with some classes missing in other depression angles. Hence, NUAAminiSAR [51] collects images of military targets at different depression angles by circular SAR in a national defense park, including 8 military vehicles and an airplane. As shown above, existing vehicle datasets lack sufficient diversity in target types, scene settings, and sensor conditions.
@@
Ship datasets - SAR ship datasets have emerged with the development of satellite SAR technology. The difference from the vehicle dataset is that the ship classification dataset usually uses a satellite platform, and the imaging angles are limited. Ship datasets can be constructed and annotated using global ship tracking systems, but label noise remains a significant challenge. OpenSARShip [26], HR4S [47], and FUSAR-Ship [27] advance the development of ship classification methods. These datasets are derived from different satellites and face difficult harbor backgrounds and class imbalance challenges. OpenSARShip collects medium-resolution images of 16 classes in harbors with Sentinel-1. HR4S builds high-resolution ship images in ports and coastal areas with GF-3 and RADARSAT-2. FUSAR-Ship contains high-resolution images under various sea backgrounds with 15 main classes and 98 subclasses of ocean targets based on GF-3. However, OpenSARShip and FUSAR-Ship have serious class imbalance and long-tailed distribution challenges. Researchers constructed various experimental settings and only used some target classes for classification tasks.
@@
Aircraft datasets - Aircraft datasets have emerged in recent years, presenting new challenges in SAR ATR. Due to their streamlined surfaces and less obvious scattering characteristics, aircraft are more difficult to classify than vehicle and ship targets. Multiangle Aircraft Target Dataset (MATD) [52] collects SAR images of two aircraft at different imaging angles with a drone strimap SAR equipment. SAR-ACD [53] contains 2 classes and 6 types of civil aircraft in 3 airports.
@@
Others - IRIS-SAR [50] is rich in aircraft, ships, and vehicle simulation types. It considers noise, clutter, and shadows in SAR image simulation. However, it does not take into account occlusion in complicated backgrounds.
@@
Object detection dataset - Beyond the classification datasets focused on individual targets, numerous SAR object detection datasets are designed to detect multiple targets. These datasets typically feature large image sizes, significant background interference, and complex targeting challenges. Examples include SARDet-100K [28] and FAIR-CSAR [29], mainly satellite-based datasets containing approximately 100,000 images with resolutions exceeding 1 meter. This limited resolution and imaging angle of satellite makes it hard to detect and classify small targets, such as vehicles, in complex scenes. Non-cooperation objectives make it difficult to label the target types and accurately control the characteristics of the target under different angles and scenes. Therefore, our dataset provides a more fine-grained SAR vehicle characterization dataset on target class and imaging conditions.
@@
Collecting target samples under different imaging conditions remains a challenging and resource-intensive endeavor. After thirty years, MSTAR remains the most used classification dataset. The evolution of SAR ATR necessitates a new large-scale dataset and benchmark in 2020s. After the above investigation, we suggest that a good SAR ATR dataset should pay attention to the following key points.
@@
Richness - A dataset should include diverse target classes, scenes, and imaging conditions to reflect real-world environments and reduce biases. It also needs to contain enough formats and metadata, such as complex data and imaging parameters.
@@
Standardization - An official dataset and benchmark are beneficial to fair research and comparisons. In addition, the dataset format and target system should be standardized to be compatible with mainstream datasets and for further expansion and scaling.
@@
Usability - A dataset should be readily accessible and user-friendly for researchers to download and utilize. Some datasets may be difficult to access or require additional SAR imaging and pre-processing steps.
@@
Therefore, we open-source a new large-scale dataset, ATRNet-STAR, with various classes and imaging conditions. This dataset is carefully labeled with extensive metadata and has an official benchmark based on 7 experimental settings and 15 comparison methods with 2 data formats. This dataset is our first step in building a new SAR target dataset system. In the future, we will expand the objects from vehicles to a wider category list to further promote SAR ATR research.